# This draft: December 17, 2025 First draft: March 5, 2025

### v1

@@ 4.2 Extracting macroeconomic state variables via autoencoder
An ablation study, presented in Internet Appendix C.5.1, further confirms the contribution of this component. Removing the autoencoder from CB-APM leads to a pronounced deterioration in predictive performance—particularly under higher $\lambda$ values—demonstrating that the learned macroeconomic embedding is essential to preserving both interpretability and accuracy. These results underscore that macroeconomic state conditioning is not a redundant extension but a core mechanism that stabilizes learning and improves out-of-sample generalization.
Finally, Appendix C.3 provides direct empirical evidence that the learned macroeconomic embeddings are economically revelatory. A two-dimensional projection of the 32-dimensional latent vectors [^10] We apply PCA to reduce the 32-dimensional latent state vectors to two dimensions only for visualization purpose. reveals a smooth temporal trajectory that coherently tracks major macroeconomic transitions, including distinct clusters corresponding to NBER-defined recessions such as the 2001 and 2008 downturns. Beyond these discrete regime shifts, the latent trajectory captures gradual cyclical and structural evolutions in the U.S. economy, reflecting shifts in growth, inflation, and monetary policy regimes. Collectively, these findings validate that the autoencoder encodes meaningful macro-financial state dynamics rather than statistical artifacts, yielding a compact and economically coherent representation that conditions firm-level return predictions within the CB-APM framework.
@@ 5.1 Cross-section of consensuses and stock returns
$R^{2}_{\text{return}}=1-\frac{\sum_{i=1}^{N}\sum_{t=1}^{T}(R_{i,t+h}-\hat{R}_{i,t+h})^{2}}{\sum_{i=1}^{N}\sum_{t=1}^{T}R_{i,t+h}^{2}},$
@@
$R^{2}_{\text{consensus}}=1-\frac{\sum_{i=1}^{N}\sum_{t=1}^{T}(C_{i,t}-\hat{C}_{i,t})^{2}}{\sum_{i=1}^{N}\sum_{t=1}^{T}C_{i,t}^{2}},$
for consensus prediction, where $N$ and $T$ denote the number of firms and time periods, respectively.
While much of the asset pricing literature emphasizes short-horizon return forecasts, sell-side analysts typically issue multi-quarter to annual forecasts. Consensus measures therefore reflect longer-term expectations about fundamentals and risk premia rather than short-term price fluctuations. Evaluating the consensus-bottleneck over horizons that align with analysts’ forecast horizons is more economically relevant than using noisy short-term intervals. Accordingly, we focus on annual return prediction, consistent with prior studies on long-horizon predictability (gu2020empirical; leippold2022machine). [^12] The results for other forecasting horizons are provided in Internet Appendix C.1
@@
To further examine robustness, Figure 7 presents results from the expanding window evaluation, comparing a naïve neural network ($\lambda=0$) to the best-performing CB-APM model ($\lambda=0.42$). The naïve network exhibits volatile performance, including negative $R^{2}$ in early periods (e.g., 2013) and isolated spikes (e.g., 2020). By contrast, CB-APM delivers consistently positive return $R^{2}$ across nearly all periods. Notably, improvements are most pronounced in periods where the naïve model already performed well (e.g., 2020–2021), indicating that CB-APM amplifies return-predictive signals in favorable regimes while mitigating underperformance in weaker ones (e.g., 2013 and 2017).
[Insert Figure 7 here]
The consensus approximation results (red bars) in the right panel further underscore CB-APM’s stability. Consensus $R^{2}$ remains high across all testing periods, demonstrating that the concept-bottleneck captures analysts’ aggregated expectations in a time-consistent manner. This stability contributes to return forecasting accuracy without overfitting to specific regimes, yielding gains across both tranquil periods (e.g., 2014–2018) and turbulent intervals such as the 2020 pandemic shock.
@@
While the out-of-sample $R^{2}$ metrics directly capture forecasting accuracy, they do not reveal how the joint loss function in equation (4) balances return prediction and consensus approximation during training. To address this, Internet Appendix C.2 provides additional evidence on the optimization dynamics of CB-APM by reporting the in-sample MSE, which demonstrates that, at short horizons, increasing $\lambda$ introduces the expected trade-off between predictive accuracy and consensus reconstruction, whereas at longer horizons the two objectives reinforce each other, yielding what we term an interpretability-accuracy amplification effect.
@@ 5.2.1 Portfolio sorts on approximated consensuses and expected returns
To further examine the pricing content embedded in CB-APM forecasts, we conduct a double-sorting exercise based on the model-implied expected returns and the analyst-based measure Analyst earnings per share (FEPS). At each month in the out-of-sample period, all stocks are first assigned to quintiles using their CB-APM-approximated FEPS levels. Within each FEPS group, stocks are then independently sorted into quintiles by their CB-APM-predicted annual returns. This procedure yields 5×5 portfolios rebalanced monthly, ensuring that both the sorting signal and subsequent return evaluation rely strictly on information available at the prediction date. For each panel, the bottom and rightmost rows report high-minus-low (H–L) spreads along the predicted-return and consensus dimensions, measuring the incremental ordering power of CB-APM forecasts conditional on FEPS.
@@ 5.2.2 Long-short portfolio performance
Portfolio turnover remains high—approximately 60% per month—which is consistent with the characteristics of complex nonlinear architectures. [^13] A formal transaction‐cost analysis based on the turnover definition in Equation (7) is provided in Internet Appendix C.4.3. The results show that the main economic conclusions are robust to proportional trading costs. This observation aligns with the findings of gu2020empirical, suggesting that neural-network-based return predictors typically produce higher turnover than linear or tree-based models due to their greater sensitivity to small shifts in cross-sectional signals. While kelly2024virtue argue that out-of-sample predictive $R^{2}$ and Sharpe ratios of characteristics-sorted portfolios may not always constitute decisive evidence of pricing relevance, the convergence of both statistical and economic measures in CB-APM suggests that its latent consensus components capture systematically priced information that conventional deep learning frameworks fail to isolate. Together, these results affirm that CB-APM’s consensus-bottleneck not only improves explanatory power but also yields tangible, risk-adjusted portfolio benefits, linking interpretability and profitability within a unified empirical asset pricing framework.
@@ 6 Dissecting Approximated Consensuses
To formalize this connection, we align our empirical strategy with standard asset pricing methodology and implement two complementary tests. First, we estimate pooled panel OLS regressions of annual stock returns on either raw analyst consensus variables or their CB-APM–inferred counterparts, thereby quantifying the incremental explanatory content gained through the consensus-bottleneck transformation. Second, we examine whether factor-mimicking portfolios—constructed from the consensus dimensions via decile sorts—span the stochastic discount factor by applying the Gibbons–Ross–Shanken (GRS) test for mean–variance efficiency (gibbons1989test). These analyses serve a dual purpose: they link the interpretability of CB-APM’s consensus layer to established empirical asset pricing tools, and they enable a direct assessment of whether the machine-inferred beliefs embody priced economic content beyond what is observable from raw analyst forecasts.
@@ 6.1 Comparative regression analysis
For the analysis, we set $h=12$ to focus on annual return predictability, thereby aligning the return horizon with prior empirical studies discussed in the preceding sections. The model is estimated on the stacked cross-section of firm-month $(i,t)$ observations to obtain a time-invariant coefficient vector $\widehat{\bm{b}}$. Inference is based on heteroskedasticity-robust covariance estimation tailored to the dependent variable’s structure. To handle an overlapping long-horizon return, we compute Driscoll–Kraay (kernel HAC) standard errors with a Bartlett kernel and an eleven-month bandwidth, which are robust to heteroskedasticity, cross-sectional dependence, and and the serial correlation induced by overlapping observations (driscoll1998consistent; newey1986simple; hodrick1992dividend).
@@ 6.2 Pricing error of test assets
The preceding analysis establishes that the CB-APM’s interpretable consensus layer captures return-relevant structure in the cross-section of individual stocks. We now examine whether these signals possess asset pricing content when evaluated through the lens of linear factor models. Specifically, we assess (i) whether the latent representations learned by CB-APM [^15] It is important to point out that the consensus representations learned by the CB-APM are not designed to approximate the span of the stochastic discount factor (SDF). Rather, the architecture learns a set of conditional expectation operators that map firm characteristics and macroeconomic conditions into consensus-like forecasts of future fundamentals. These latent expectations summarize belief-based or information-based heterogeneity, not compensated sources of systematic risk. Consequently, the CB-APM should not be expected to replicate the factor structure implicit in linear SDF models; instead, its consensus layer provides an economically interpretable decomposition of expected returns that is complementary to—rather than a substitute for—the traditional factor space. can serve as risk factors capable of pricing standard benchmark portfolios, and (ii) whether traditional factor models can price the return patterns implied by the CB-APM’s predictions and consensus-based characteristics. To do so, we employ the multivariate Gibbons–Ross–Shanken (GRS) test (gibbons1989test), which jointly evaluates whether the intercepts ($\bm{\alpha}$) in time-series regressions are statistically different from zero.
@@
Tables 6LABEL:tab:grs-consensus present a comprehensive set of in-sample Gibbons–Ross–Shanken (GRS) tests evaluating the pricing performance of the CB-APM relative to conventional factor models. Across all tests, the GRS $F$-statistic assesses the joint null hypothesis that all pricing errors ($\bm{\alpha}$) are zero, such that lower $F$-statistics and higher $p$-values indicate superior mean–variance efficiency. The accompanying mean absolute and root-mean-squared (RMS) alphas summarize the magnitude of mispricing across the corresponding test assets.
@@
Panels A–C of Table 6 evaluate whether the CB-APM’s consensus-based factor-mimicking portfolios can price the returns of the Fama–French 25 size–book-to-market portfolios, the 25 size–momentum portfolios, and the 30 industry portfolios. Across these benchmarks, the CB-APM factors deliver GRS statistics that are broadly comparable to those of standard models, but they remain somewhat higher than the FF5 and FF6 specifications. Mean and RMS pricing errors are likewise modest yet consistently larger than those generated by traditional factor structures. These results indicate that the consensus-based factors span meaningful components of systematic return variation, but not the full set captured by benchmark style factors. This is consistent with evidence that only a limited number of characteristic-based factors are strongly priced in the cross-section, while many signals are redundant or weakly informative (kozak2020shrinking). Within this environment, the CB-APM factors behave as an additional block of characteristic-sorted portfolios that contributes incremental explanatory variation without supplanting the canonical factor structure.
@@
The three sets of GRS tests reveal how the CB-APM relates to the traditional factor space. First, consensus-based factor-mimicking portfolios do not fully price the classic benchmark portfolios, which indicates that the latent consensus dimensions do not function as close substitutes for the core priced factors. Second, the ability of traditional factor models to price CB-APM–generated portfolios deteriorates as the consensus-bottleneck becomes more stringent, implying that the model’s predictive signals progressively move outside the span of standard linear characteristics. This behavior is consistent with the broader view that, while the priced dimension of the stochastic discount factor is relatively low, flexible methods can uncover structured forms of return heterogeneity that improve portfolio efficiency (cong2025growing) without reproducing the canonical factors directly. Third, portfolios sorted on individual consensus dimensions exhibit moderate but non-negligible pricing errors, suggesting that the learned signals contain meaningful information about expected returns but do not themselves constitute a new stand-alone factor system.
This finding crucially aligns with the emerging methodological consensus that traditional characteristic–based sorting procedures fundamentally fail to capture the full mean–variance efficient (MVE) frontier due to their neglect of nonlinearity and asymmetric characteristic interactions. Recent goal-oriented machine learning approaches—most notably the Panel Tree (P-Tree) framework (cong2025growing) and the Asset Pricing Tree (AP-Tree) framework (bryzgalova2025forest)—demonstrate that test assets constructed by explicitly optimizing for SDF spanning or MVE efficiency are substantially harder to price with conventional factor models, often yielding extremely high GRS statistics. In this sense, the behavior observed in Table 7 echoes the insight that once test assets begin to reflect structured, state-dependent return heterogeneity, linear factor structures fail sharply.
@@ 7 Conclusion
This study introduces the Consensus-Bottleneck Asset Pricing Model (CB-APM), a novel framework that integrates interpretable deep learning with empirical asset pricing. By embedding a concept-bottleneck architecture into a neural network, CB-APM not only achieves state-of-the-art predictive accuracy in cross-sectional stock return forecasts but also provides transparent insights into the role of analysts’ consensus in shaping risk premiums. Our empirical results demonstrate that interpretability and performance are not inherently conflicting. CB-APM outperforms conventional deep learning benchmarks in long-horizon forecasts while preserving a clear, economically grounded structure. By linking machine learning’s predictive capabilities with the theoretical underpinnings of financial economics, and by demonstrating that interpretable deep learning can yield both statistical and economic validity, this work offers a blueprint for building models that are both high-performing and aligned with established asset pricing principles.
@@
Table 3: Realized monthly returns of out-of-sample double-sorted portfolios across $\lambda$. Each panel reports mean monthly realized returns (in percentage points) for monthly rebalanced 5×5 portfolios sorted by the approximated Analyst earning per share (E[FEPS]E[{FEPS}], rows) and predicted annual returns ($E[R]$, columns), independently. H–L denotes the high-minus-low spread across the corresponding dimension.
@@
$E_{t}[R_{i,t+h}]$
$E_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$E_{t}[R_{i,t+h}]$
$E_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$E_{t}[R_{i,t+h}]$
$E_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$E_{t}[R_{i,t+h}]$
$E_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$E_{t}[R_{i,t+h}]$
$E_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$E_{t}[R_{i,t+h}]$
$E_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$E_{t}[R_{i,t+h}]$
$E_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$E_{t}[R_{i,t+h}]$
$E_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$E_{t}[R_{i,t+h}]$
$E_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$E_{t}[R_{i,t+h}]$
$E_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
This table reports in-sample Gibbons–Ross–Shanken (GRS) tests of mean–variance efficiency for competing factor models.The test assets include (i) the 25 size–book-to-market portfolios, (ii) the 25 size–momentum portfolios, and (iii) the 30 value-weighted industry portfolios. CB-APM factors are constructed by sorting stocks into deciles on each consensus dimension and taking the value-weighted return spread between the top and bottom deciles, yielding tradable factor-mimicking portfolios.Different values of $\lambda$ correspond to CB-APM models trained under varying strengths of the consensus-bottleneck.Each panel reports the GRS $F$-statistic, $p$-value, mean absolute and RMS pricing errors (monthly and annualized), and the number of factors ($K$). All statistics use monthly excess returns.
@@
This table reports in-sample Gibbons–Ross–Shanken (GRS) tests applied to decile portfolios constructed from CB-APM predicted return scores. Each $\lambda$ corresponds to a distinct CB-APM specification that generates the test assets. For each $\lambda$, we report the GRS $F$-statistic, its $p$-value, and mean absolute and root-mean-squared pricing errors (monthly and annualized). All models are estimated using monthly excess returns.
@@
Table 8: GRS tests for traditional factor models applied to portfolios sorted on CB-APM approximated consensus signals. This table reports Gibbons–Ross–Shanken (GRS) test statistics for value-weighted decile portfolios formed on each dimension of the CB-APM’s approximated consensus at $\lambda=1.0$. For each consensus dimension, portfolio returns are regressed on the CAPM, Fama–French three-factor model, Carhart four-factor model, Fama–French five-factor model, and Fama–French six-factor model. Reported are the GRS $F$-statistic, corresponding $p$-value, and mean absolute and root-mean-squared pricing errors, shown in both monthly and annualized units.
@@
Figure 1: Disentangled representations of neural network.
Consensus Module $f(\phi)$ Prediction Module $g(\theta)$ Predictors $I^{f}_{i,t},I^{m}_{t}$ Consensus $\hat{C}_{i,t}$Excess Return $E_{t}[R_{i,t+h}]$$L_{R}(\phi,\theta)$$\lambda\,L_{c}(\phi)$$L$
Figure 2: Architecture of the Consensus-Bottleneck Asset Pricing Model (CB-APM).
The model is composed of two modules, the consensus module $f(\phi)$ (left) and the prediction module $g(\theta)$ (right). The consensus module compresses firm-specific predictors $I^{f}_{i,t}$ and macroeconomic variables $I^{m}_{t}$ into a lower-dimensional consensus vector $\hat{C}_{i,t}$ through a feedforward neural network. This bottleneck enforces interpretability by design, as each coordinate of $\hat{C}_{i,t}$ is treated as a consensus concept. The prediction module then maps these consensus variables into expected excess returns $E_{t}[R_{i,t+h}]$ using a linear layer. The return loss $L_{R}(\phi,\theta)$ and the consensus loss $L_{c}(\phi)$ are optimized jointly using the weighted sum $L=\lambda L_{c}+L_{R}$, ensuring that the consensus layer is both predictive of returns and interpretable.
@@
Figure 7: Out-of-sample $R^{2}$ by testing period under expanding window evaluation.
This figure reports monthly $R^{2}$ of annual stock return and consensus prediction by testing period, based on an expanding-window evaluation. The left panel presents results for a naïve neural network without consensus learning ($\lambda=0$), while the right panel shows results for the best-performing model ($\lambda=0.42$).

### v2

@@ 3.2 Extracting macroeconomic state variables via autoencoder
An ablation study, presented in Internet Appendix D.5.1, further confirms the contribution of this component. Removing the autoencoder from CB-APM leads to a pronounced deterioration in predictive performance—particularly under higher $\lambda$ values—demonstrating that the learned macroeconomic embedding is essential to preserving both interpretability and accuracy. These results underscore that macroeconomic state conditioning is not a redundant extension but a core mechanism that stabilizes learning and improves out-of-sample generalization.
Finally, Appendix D.3 provides direct empirical evidence that the learned macroeconomic embeddings are economically revelatory. A two-dimensional projection of the 32-dimensional latent vectors [^9] We apply PCA to reduce the 32-dimensional latent state vectors to two dimensions only for visualization purpose. reveals a smooth temporal trajectory that coherently tracks major macroeconomic transitions, including distinct clusters corresponding to National Bureau of Economic Research(NBER)-defined recessions such as the 2001 and 2008 downturns. Beyond these discrete regime shifts, the latent trajectory captures gradual cyclical and structural evolutions in the U.S. economy, reflecting shifts in growth, inflation, and monetary policy regimes. Collectively, these findings validate that the autoencoder encodes meaningful macro-financial state dynamics rather than statistical artifacts, yielding a compact and economically coherent representation that conditions firm-level return predictions within the CB-APM framework.
@@ 4.1 Cross-section of consensuses and stock returns
$R^{2}_{\text{return}}=1-\frac{\sum_{i=1}^{N}\sum_{t=1}^{T}\bigl(R_{i,t+h}-\hat{R}_{i,t+h}\bigr)^{2}}{\sum_{i=1}^{N}\sum_{t=1}^{T}R_{i,t+h}^{2}},$
@@
$R^{2}_{\text{consensus}}=1-\frac{\sum_{i=1}^{N}\sum_{t=1}^{T}\bigl(C_{i,t}-\hat{C}_{i,t}\bigr)^{2}}{\sum_{i=1}^{N}\sum_{t=1}^{T}C_{i,t}^{2}},$
for consensus approximation, where $N$ and $T$ denote the number of firms and time periods, respectively.
While much of the asset pricing literature emphasizes short-horizon return forecasts, sell-side analysts typically issue multi-quarter to annual forecasts. Consensus measures therefore reflect longer-term expectations about fundamentals and risk premia rather than short-term price fluctuations. Evaluating the consensus-bottleneck over horizons that align with analysts’ forecast horizons is more economically relevant than using noisy short-term intervals. Accordingly, we focus on annual return prediction, consistent with prior studies on long-horizon predictability (gu2020empirical; leippold2022machine). [^11] The results for other forecasting horizons are provided in Internet Appendix D.1
@@
While the out-of-sample $R^{2}$ metrics directly capture forecasting accuracy, they do not reveal how the joint loss function in equation (4) balances return prediction and consensus approximation during training. To address this, Internet Appendix D.2 provides additional evidence on the optimization dynamics of CB-APM by reporting the in-sample MSE, which demonstrates that, at short horizons, increasing $\lambda$ introduces the expected trade-off between predictive accuracy and consensus reconstruction, whereas at longer horizons the two objectives reinforce each other, yielding what we term an interpretability-accuracy amplification effect.
@@ 4.2.1 Portfolio sorts on approximated consensuses and expected returns
To further examine the pricing content embedded in CB-APM forecasts, we conduct a double-sorting exercise based on the model-implied expected returns and the analyst-based measure FEPS. At each month in the out-of-sample period, all stocks are first assigned to quintiles using their CB-APM-approximated FEPS levels. Within each FEPS group, stocks are then independently sorted into quintiles by their CB-APM-predicted annual returns. This procedure yields 5×5 portfolios rebalanced monthly, ensuring that both the sorting signal and subsequent return evaluation rely strictly on information available at the prediction date. For each panel, the bottom and rightmost rows report high-minus-low (H–L) spreads along the predicted-return and consensus dimensions, measuring the incremental ordering power of CB-APM forecasts conditional on FEPS.
@@ 4.2.2 Long-short portfolio performance
Portfolio turnover remains high—approximately 60% per month—which is consistent with the characteristics of complex nonlinear architectures. [^12] A formal transaction‐cost analysis based on the turnover definition in Equation (7) is provided in Internet Appendix D.4.3. The results show that the main economic conclusions are robust to proportional trading costs. This observation aligns with the findings of gu2020empirical, suggesting that neural-network-based return predictors typically produce higher turnover than linear or tree-based models due to their greater sensitivity to small shifts in cross-sectional signals. While kelly2024virtue argue that out-of-sample predictive $R^{2}$ and Sharpe ratios of characteristics-sorted portfolios may not always constitute decisive evidence of pricing relevance, the convergence of both statistical and economic measures in CB-APM suggests that its latent consensus components capture systematically priced information that conventional deep learning frameworks fail to isolate. Together, these results affirm that CB-APM’s consensus-bottleneck not only improves explanatory power but also yields tangible, risk-adjusted portfolio benefits, linking interpretability and profitability within a unified empirical asset pricing framework.
@@ 5 Dissecting Approximated Consensuses
To formalize this connection, we align our empirical strategy with standard asset pricing methodology and implement two complementary tests. First, we estimate pooled panel OLS regressions of annual stock returns on either raw analyst consensus variables or their CB-APM–inferred counterparts, thereby quantifying the incremental explanatory content gained through the consensus-bottleneck transformation. Second, we examine whether factor-mimicking portfolios—constructed from the consensus dimensions via decile sorts—span the SDF by applying the GRS test for mean–variance efficiency (gibbons1989test). These analyses serve a dual purpose: they link the interpretability of CB-APM’s consensus layer to established empirical asset pricing tools, and they enable a direct assessment of whether the machine-inferred beliefs embody priced economic content beyond what is observable from raw analyst forecasts.
@@ 5.1 Comparative regression analysis
For the analysis, we set $h=12$ to focus on annual return predictability, thereby aligning the return horizon with prior empirical studies discussed in the preceding sections. The model is estimated on the stacked cross-section of firm-month $(i,t)$ observations to obtain a time-invariant coefficient vector $\widehat{\bm{b}}$. Inference is based on heteroskedasticity-robust covariance estimation tailored to the dependent variable’s structure. To handle an overlapping long-horizon return, we compute Driscoll–Kraay (kernel HAC) standard errors with a Bartlett kernel and an eleven-month bandwidth, which are robust to heteroskedasticity, cross-sectional dependence, and the serial correlation induced by overlapping observations (driscoll1998consistent; newey1986simple; hodrick1992dividend).
@@ 5.2 Pricing error of test assets
The preceding analysis establishes that the CB-APM’s interpretable consensus layer captures return-relevant structure in the cross-section of individual stocks. We now examine whether these signals possess asset pricing content when evaluated through the lens of linear factor models. Specifically, we assess (i) whether the latent representations learned by CB-APM [^14] It is important to point out that the consensus representations learned by the CB-APM are not designed to approximate the span of the SDF. Rather, the architecture learns a set of conditional expectation operators that map firm characteristics and macroeconomic conditions into consensus-like forecasts of future fundamentals. These latent expectations summarize belief-based or information-based heterogeneity, not compensated sources of systematic risk. Consequently, the CB-APM should not be expected to replicate the factor structure implicit in linear SDF models; instead, its consensus layer provides an economically interpretable decomposition of expected returns that is complementary to—rather than a substitute for—the traditional factor space. can serve as risk factors capable of pricing standard benchmark portfolios, and (ii) whether traditional factor models can price the return patterns implied by the CB-APM’s predictions and consensus-based characteristics. To do so, we employ the multivariate GRS test (gibbons1989test), which jointly evaluates whether the intercepts ($\bm{\alpha}$) in time-series regressions are statistically different from zero.
@@
Tables 6LABEL:tab:grs-consensus present a comprehensive set of in-sample GRS tests evaluating the pricing performance of the CB-APM relative to conventional factor models. Across all tests, the GRS $F$-statistic assesses the joint null hypothesis that all pricing errors ($\bm{\alpha}$) are zero, such that lower $F$-statistics and higher $p$-values indicate superior mean–variance efficiency. The accompanying mean absolute and root-mean-squared alphas summarize the magnitude of mispricing across the corresponding test assets.
@@
Panels A–C of Table 6 evaluate whether the CB-APM’s consensus-based factor-mimicking portfolios can price the returns of the Fama–French 25 size–book-to-market portfolios, the 25 size–momentum portfolios, and the 30 industry portfolios. Across these benchmarks, the CB-APM factors deliver GRS statistics that are broadly comparable to those of standard models, but they remain somewhat higher than the Fama–French five-factor model and Fama–French six-factor model specifications. Mean and RMS pricing errors are likewise modest yet consistently larger than those generated by traditional factor structures. These results indicate that the consensus-based factors span meaningful components of systematic return variation, but not the full set captured by benchmark style factors. This is consistent with evidence that only a limited number of characteristic-based factors are strongly priced in the cross-section, while many signals are redundant or weakly informative (kozak2020shrinking). Within this environment, the CB-APM factors behave as an additional block of characteristic-sorted portfolios that contributes incremental explanatory variation without supplanting the canonical factor structure.
@@
The three sets of GRS tests reveal how the CB-APM relates to the traditional factor space. First, consensus-based factor-mimicking portfolios do not fully price the classic benchmark portfolios, which indicates that the latent consensus dimensions do not function as close substitutes for the core priced factors. Second, the ability of traditional factor models to price CB-APM–generated portfolios deteriorates as the consensus-bottleneck becomes more stringent, implying that the model’s predictive signals progressively move outside the span of standard linear characteristics. This behavior is consistent with the broader view that, while the priced dimension of the SDF is relatively low, flexible methods can uncover structured forms of return heterogeneity that improve portfolio efficiency (cong2025growing) without reproducing the canonical factors directly. Third, portfolios sorted on individual consensus dimensions exhibit moderate but non-negligible pricing errors, suggesting that the learned signals contain meaningful information about expected returns but do not themselves constitute a new standalone factor system.
This finding crucially aligns with the emerging methodological consensus that traditional characteristic–based sorting procedures fundamentally fail to capture the full mean–variance efficient (MVE) frontier due to their neglect of nonlinearity and asymmetric characteristic interactions. Recent goal-oriented machine learning approaches—most notably the Panel Tree (P-Tree) framework (cong2025growing) and the Asset Pricing Tree (AP-Tree) framework (bryzgalova2025forest)—demonstrate that test assets constructed by explicitly optimizing for SDF spanning or mean–variance efficiency are substantially harder to price with conventional factor models, often yielding extremely high GRS statistics. In this sense, the behavior observed in Table 7 echoes the insight that once test assets begin to reflect structured, state-dependent return heterogeneity, linear factor structures fail sharply.
@@ 6 Conclusion
This study introduces the CB-APM, a novel framework that integrates interpretable deep learning with empirical asset pricing. By embedding a concept-bottleneck architecture into a neural network, CB-APM not only achieves state-of-the-art predictive accuracy in cross-sectional stock return forecasts but also provides transparent insights into the role of analysts’ consensus in shaping risk premiums. Our empirical results demonstrate that interpretability and performance are not inherently conflicting. CB-APM outperforms conventional deep learning benchmarks in long-horizon forecasts while preserving a clear, economically grounded structure. By linking machine learning’s predictive capabilities with the theoretical underpinnings of financial economics, and by demonstrating that interpretable deep learning can yield both statistical and economic validity, this work offers a blueprint for building models that are both high-performing and aligned with established asset pricing principles.
@@
Table 3: Realized monthly returns of out-of-sample double-sorted portfolios across $\lambda$. Each panel reports mean monthly realized returns (in percentage points) for monthly rebalanced 5×5 portfolios sorted by the approximated Analyst earning per share ($\mathbb{E}\left[{FEPS}\right]$, rows) and predicted annual returns ($\mathbb{E}\left[R\right]$, columns), independently. H–L denotes the high-minus-low spread across the corresponding dimension.
@@
$\mathbb{E}_{t}[R_{i,t+h}]$
$\mathbb{E}_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$\mathbb{E}_{t}[R_{i,t+h}]$
$\mathbb{E}_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$\mathbb{E}_{t}[R_{i,t+h}]$
$\mathbb{E}_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$\mathbb{E}_{t}[R_{i,t+h}]$
$\mathbb{E}_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$\mathbb{E}_{t}[R_{i,t+h}]$
$\mathbb{E}_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$\mathbb{E}_{t}[R_{i,t+h}]$
$\mathbb{E}_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$\mathbb{E}_{t}[R_{i,t+h}]$
$\mathbb{E}_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$\mathbb{E}_{t}[R_{i,t+h}]$
$\mathbb{E}_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$\mathbb{E}_{t}[R_{i,t+h}]$
$\mathbb{E}_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
$\mathbb{E}_{t}[R_{i,t+h}]$
$\mathbb{E}_{t}[{FEPS}_{i,t}]$ Low 22 33 44 High HL
@@
This table reports in-sample GRS tests of mean–variance efficiency for competing factor models.The test assets include (i) the 25 size–book-to-market portfolios, (ii) the 25 size–momentum portfolios, and (iii) the 30 value-weighted industry portfolios. CB-APM factors are constructed by sorting stocks into deciles on each consensus dimension and taking the value-weighted return spread between the top and bottom deciles, yielding tradable factor-mimicking portfolios.Different values of $\lambda$ correspond to CB-APM models trained under varying strengths of the consensus-bottleneck.Each panel reports the GRS $F$-statistic, $p$-value, mean absolute and RMS pricing errors (monthly and annualized), and the number of factors ($K$). All statistics use monthly excess returns.
@@
This table reports in-sample GRS tests applied to decile portfolios constructed from CB-APM predicted return scores. Each $\lambda$ corresponds to a distinct CB-APM specification that generates the test assets. For each $\lambda$, we report the GRS $F$-statistic, its $p$-value, and mean absolute and root-mean-squared pricing errors (monthly and annualized). All models are estimated using monthly excess returns.
@@
Table 8: GRS tests for traditional factor models applied to portfolios sorted on CB-APM approximated consensus signals. This table reports GRS test statistics for value-weighted decile portfolios formed on each dimension of the CB-APM’s approximated consensus at $\lambda=1.0$. For each consensus dimension, portfolio returns are regressed on the CAPM, Fama–French three-factor model, Carhart four-factor model, Fama–French five-factor model, and Fama–French six-factor model. Reported are the GRS $F$-statistic, corresponding $p$-value, and mean absolute and root-mean-squared pricing errors, shown in both monthly and annualized units.
@@
Consensus Module $f(\phi)$ Prediction Module $g(\theta)$ Predictors $\mathcal{I}^{f}_{i,t},\mathcal{I}^{m}_{t}$ Consensus $\hat{C}_{i,t}$Excess Return $\mathbb{E}_{t}\left[R_{i,t+h}\right]$$\mathcal{L}_{R}(\phi,\theta)$$\lambda\,\langle\textbf{1},\mathcal{L}_{C}(\phi)\rangle\mathcal{L}$
Figure 1: Architecture of the CB-APM.
The model is composed of two modules, the consensus module $f(\phi)$ (left) and the prediction module $g(\theta)$ (right). The consensus module compresses firm-specific predictors $\mathcal{I}^{f}_{i,t}$ and macroeconomic variables $\mathcal{I}^{m}_{t}$ into a lower-dimensional consensus vector $\hat{C}_{i,t}$ through a feedforward neural network. This bottleneck enforces interpretability by design, as each coordinate of $\hat{C}_{i,t}$ is treated as a consensus concept. The prediction module then maps these consensus variables into expected excess returns $\mathbb{E}_{t}\left[R_{i,t+h}\right]$ using a linear layer. The return loss $\mathcal{L}_{R}(\phi,\theta)$ and the consensus loss $\mathcal{L}_{C}(\phi)$ are optimized jointly using the weighted sum $\mathcal{L}=\mathcal{L}_{R}+\lambda\,\langle\textbf{1},\mathcal{L}_{C}\rangle$, ensuring that the consensus layer is both predictive of returns and interpretable.